U=(9,2) v=(-5,-2) 2u-3v=?
Answer by jim_thompson5910(35256)

u = <9, 2>
2*u = 2*<9, 2>
2*u = <2*9, 2*2>
2*u = <18, 4>


v = <-5, -2>
-3*v = -3*<-5, -2>
-3*v = <-3*(-5), -3*(-2)>
-3*v = <15, 6>


-----------------------------------------

2u - 3v = 2u + (-3v)
2u - 3v = (<18, 4>) + (<15, 6>)
2u - 3v = <18, 4> + <15, 6>
2u - 3v = <18+15, 4+6>
2u - 3v = <33, 10>